Monochromatic light of wavelength 500 nm is incident on a narrow slit. On a screen 2.01 m away, the distance between the second diffraction minimum and the central maximum is 1.85 cm. (a) Calculate the angle of diffraction ? of the second minimum. (b) Find the width of the slit.
Answer #1

lambda=500 nm


distace between the slit and screen, D=2.01 m


y2=1.85cm

a)

use,


tan(theta)=y2/D


=1.85*10^-2/(2.01)

===> angle of diffraction, theta=0.527 degrees

b)


d*sin(theta)=m*lambda


d*sin(0.527)=2*453*10^-9

===> slit width, d=9.85*10^-5 m or 98.5 um